# Fenchel-Young Variational Inference

### v2

@@ 3.4 Gradient Backpropagation
In two of the experiments discussed below, we optimize the FYELBO in an amortized inference setting [24], where the variational distribution is parametrized by an encoder network that takes $x$ as input, as in variational autoencoders (VAE) [33]. We denote the resulting variational distribution by $q_{\phi}(z|x)$, where $\phi$ are the inference network parameters. Instead of minimizing (14) through alternating minimization, we use end-to-end gradient backpropagation combined with a reparametrization trick to update the encoder parameters $\phi$, the decoder parameters $\theta$, and, optionally, any parameters associated with $\eta$ (our FYVAEs use standard fixed priors). It is also obviously possible to use a score function estimator instead of reparametrization.
@@ 4 Experiments
We illustrate FYVI through in several applications. The first is GMM-based clustering, where we recover the standard and hard versions of EM for particular choices of regularizer, in addition to a new sparse [^4] This sparse variant of EM is a different instance of FYVI than those proposed by Dukkipati et al. [21] and Inoue and Shinoda [29], which are standard mixtures of $(2-\rho)$-Gaussians. EM. We then use FYVI in $\beta$-VAEs with finite posterior support and apply them to MNIST image data with both sparse and standard multivariate Bernoulli observation models, and to Fashion MNIST data with sparse and standard Categorical observation models. Finally, we combine a Gaussian prior with sparse observation models to model a sparse empirical distribution in a neural variational document model.
@@
Our focus is not computational scalability. Depending on the particular model, parameter $\rho$ in Tsallis $\rho$-negentropy, and parametrization, the computational costs will be different. Computational cost depends on model-specific choices, rather than the generalization of standard VI to FYVI. Tradeoffs between FY and KL regularization similarly depend on model specification; there is no one general comparison. In most of our experiments, computational tradeoffs between FY regularizers and KL regularizers are negligible. There are efficient algorithms for computing sparsemax ($\rho=2$) and $1.5$-entmax; otherwise computing softmax is faster. We use reparametrization tricks in all the VAE experiments: the classical Gaussian one [9], and a $(2-\rho)$-Gaussian reparametrization [39]. We keep computational resources as close as possible between the compared models.
4.1 FYEM for GMM
To instantiate the FYEM algorithm described above to estimate a Gaussian mixture model (GMM), we set $\mathcal{L}_{\theta}(x_{i},z_{i})=-\log\mathcal{N}(x_{i};\mu_{z_{i}},\Sigma_{z_{i}})$ as Gaussian negative log-likelihoods. Setting $\Omega=\Omega_{1}$ recovers the standard EM algorithm for GMM estimation. We also consider other regularizers, $\Omega_{\rho}$ with $p>1$, which yield a new adaptively sparse EM algorithm for GMMs. In Appendix B we review of the baseline EM and hard EM algorithms compared in Figure 1, give details of how the clusters are generated and evaluated in Table 1, and an analysis of the sparsity of E-steps. Figure 2 illustrates how average sparsity and the unuspervised clustering evaluation metrics we use vary with $\rho$.
Figure 1: Comparison of the standard, sparse, and hard versions of EM for GMM-based clustering. Data points are colored by their true labels, with x markers indicating outliers. Black ellipses are level curves of the estimated GMM components, while colored ellipses denote the ground-truth.
Setting $\Omega=\Omega_{\rho}$, the Tsallis negentropy (10) with $\rho>1$, yields the new sparse version of (15):
${\hat{q}_{i}^{\mbox{\scriptsize new}}}=\rho\mbox{-entmax}\Bigl{(}\frac{\pi_{1}^{\rho-1}}{\rho-1}+\log\bigl{(}\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1})\bigr{)},...,\frac{\pi_{K}^{\rho-1}}{\rho-1}+\log\bigl{(}\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K}\bigr{)}\Bigr{)}.$
For $\rho>1$, the output of $\rho\mbox{-entamax}$ can have both zero and non-zero components. When $\rho=1$, Tsallis $\rho$-negentropy becomes the Shannon nentropy and we recover the standard EM algorithm, where all the components of $\hat{q}_{i}^{\mbox{\scriptsize new}}$ are strictly positive. The novel feature of sparse EM is the possible presence of zeros in $\hat{q}_{i}$, meaning that sample $x_{i}$ will not contribute to the update of the parameter estimates of the corresponding components. In standard EM all samples affect the parameter estimates of all the components. In hard EM, each sample only affects the parameters of the “closest" component, as in the case of $K$-means clustering (with the important difference that $K$-means clustering does not estimate component covariances).
@@ 4.1 FYEM for GMM
Figure 2: We vary $\rho\in\{0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0\}$ and plot the resulting adjusted mutual information, adjusted Rand index, and silhouette score. If $\rho$ is close to 1, the value of metric is close to its value with classical EM. We quantify the E-step sparsity by counting the number of clusters assigned 0 mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned 0 proportion is three. The shaded regions represent standard deviations from the mean values across 5 seeds.
@@
Table 1: Clustering quality metrics (higher is better) for GMM-based clustering using the standard, hard, and sparse (ours) versions of EM; $\rho=2$ in sparse EM. Values are averaged over 5 random seeds. Priors are learned according to Eq. (16).
E-Step Adjusted MI Adjusted RI Silhouette Sc.
Standard EM .606 $\pm$ .012 .531 $\pm$ .005 .345 $\pm$ .032
Hard EM .537 $\pm$ .088 .348 $\pm$ .043 .207 $\pm$ .128
Sparse EM .636 $\pm$ .017 .476 $\pm$ .037 .393 $\pm$ .041
@@ 4.2 Fenchel-Young $\beta$-VAE
In Kingma and Welling [33]’s classic VAE a Gaussian variational posterior and a $\mathcal{N}(0,I)$ prior are combined with a product of independent Bernoulli’s as observation model, representing the pixels of MNIST digit images. The standard VI baseline $\beta$-VAE objective is included in in Appendix C. We experiment with truncated parabola $(2-\rho)$-Gaussian [^5] Martins et al. [39] call this family $\beta$-Gaussian. finite-support variational posteriors, comparing the baseline product of Bernoulli’s decoder with a product of binary $\Psi_{2}$ deformed exponential family distributions. The FYELBO objective is
$\sum\limits_{i=1}^{N}\underset{Z\sim q_{i}}{\mathbb{E}}\left[L_{\Psi_{\rho^{\prime}}}(\theta(Z);x_{i})\right]+\beta L_{\Omega_{\rho}}(\eta_{\mathcal{N}(0,\textnormal{I})};q_{i}),$
@@
Table 2: VAE $\ell_{1}$ reconstruction error on MNIST and Fashion MNIST datasets.
Observation Latent Variable Mnist FashionMnist
Product of independent Bernoulli’s, $\rho^{\prime}=1$ Gaussian, $\rho=1$ 13.272 34.192
Biweight, $\rho=3/2$ 12.061 32.690
Epanechnikov, $\rho=2$ 12.111 32.482
Product of $\Psi_{2}$ sparse binary distributions, $\rho^{\prime}=2$ Gaussian, $\rho=1$ 12.610 32.986
Biweight, $\rho=3/2$ 10.326 25.985
Epanechnikov, $\rho=2$ 9.183 26.186
In Table2 we report the results of computationally efficient combinations of $\Psi$ and $\Omega$ (i.e. strictly comparable with standard VI VAEs) in VAEs applied to MNIST and FashionMNIST. It is noteworthy that sparsity in the encoder and in the encoder separately contribute to improved performance. We report average $l_{1}$ reconstruction error as a means of comparing the different models, since the FY ELBO values themselves are not comparable. In all cases we tune $\beta$ by cross-validation to prevent the posterior from collapsing to the prior.
In Appendix D we report improved $l_{1}$ reconstruction error in a small document model VAE, using sparsity only in the observation model. This can be seen as a generalized Bayesian treatment of sparse observation models, falling under the cases of generalized VI described by Bissiri et al. [7], Khan and Rue [32], Knoblauch et al. [34]. In preliminary experiments with sparse posteriors we did not observe improvements over the baseline NVDM of Miao et al. [40]. This is intuitively reasonable: While the empirical distribution over words is sparse, one latent Multivariate Gaussian associated with each Multinomial (or $\Omega_{\rho}$-sparse) observation through amortized variational inference does not lead to improved reconstruction if its domain is truncated. In the image VAE data, similar patterns of pixels are 0 across images, e.g. on the boarders, while in the document model VAEs there is no word of the vocabulary that ought to be assigned no probability mass in all “topic dimensions”.
@@ 5 Conclusion
We introduce Fenchel-Young variational inference (FYVI), expanding the class of variational posteriors to the $(2-\rho)$-deformed exponential family. FYVI exploits a general class of entropic regularizers to obtain sparse discrete and continuous posterior distributions, and distinctly extends classical Bayesian inference compared to previous generalized Bayesian and variational inference proposals like those of Bissiri et al. [7], Knoblauch et al. [34], and Daudel et al. [18]. We focus on cases where the entropic index $\rho$ is greater than $1$, which leads to sparse posteriors previously unexplored in the related literature. Note the nature of sparsity of $(2-\rho)$-exponential family posteriors is qualitatively distinct from the sparse MAP estimates that are obtained through sparsity-inducing priors (§3). We define and lower bound the FY variational free energy (FYVFE), and describe EM-style and amortized VI algorithms for estimating arbitrary generative FYVI models. We derive and experiment with a sparse EM algorithm for Gaussian mixtures, demonstrating better clustering quality than classical and hard versions of EM on synthetic data with overlapping clusters and outliers. In our image FYVAE experiments with continuous finite support posteriors, we obtain the best performance when combining a latent $(2-\rho)$-Gaussian posterior with a sparse support observation model. We also show how to treat sparse observation models from a generalized Bayes perspective.
@@
FYVI can be directly extended to structured latent variables by using SparseMAP [46] posteriors, intuitively sparse analogues to conditional random fields. In future work we plan to extend our analysis and experiments to the non-trivial case of heavy tailed posteriors ($\rho$ < 1). Further, FYVI is a necessary step toward characterizing conjugate FY families and generalized canonical gradient descent, to account for the geometry of the space of deformed exponential family distributions, and to extend information-theoretic generalization bounds that have been studied in the case of SGD [45, 14] to locally reparametrization-robust optimization methods. Another interesting direction is to connect FYVI with variational Rényi objectives [17, 18]. The recently popular $\lambda$-deformation [60, 57, 26], which aims to treat Rényi and Tsallis entropies in a unified way, may provide an illuminating bridge among alternative generalized VI perspectives. Finally, connections with PAC-Bayes analyses of VI like Alquier et al. [3] are also an interesting prospect.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A $\rho$-MaxEnt and relation with Amari’s $\alpha$- and $q$-divergences
$\rho$-deformation definitions.
@@ Appendix A $\rho$-MaxEnt and relation with Amari’s $\alpha$- and $q$-divergences
Defining the Tsallis entropy relies on the definitions of the $\rho$-deformed logarithmic and exponential functions for an entropic index $\rho\in\mathbb{R}\setminus\{1\}$ — this index is called $q$ by Tsallis [53], Amari and Ohara [4], Naudts [43], and $\alpha$ by Martins et al. [39]:
$\log_{\rho}(x):=\frac{x^{1-\rho}-1}{1-\rho},\quad x>0,\quad\exp_{\rho}(x):=[1+(1-\rho)x]_{+}^{\frac{1}{1-\rho}},\quad x\in\mathbb{R}$ (18)
where $\cdot]_{+}:=\max\{\cdot,0\}$, which in the limit $\rho\rightarrow 1$ converge to $\log(x)$ and $\exp(x)$. Note the following relations:
@@
$L_{\Omega}$ and related divergences.
Eq. (10) above defines the Tsallis $\rho$-negentropy scaled by $\frac{1}{\rho}$ to preserve consistency with Martins et al. [39] $\Omega_{\rho}(q)=-\frac{1}{\rho}S_{\rho}(q)$ where $S_{\rho}(q)=\frac{1}{1-\rho}\mathbb{E}_{q(z)}\left[q(z)^{\rho-1}-1\right]$ is the Tsallis entropy. Additionally, noting that $\log_{\rho}(\frac{1}{q})+\log_{2-\rho}q=0$
@@
The Tsallis $\rho$-relative entropy, given by Amari and Ohara [4, Eq. 39] is a scalar multiple of Amari and Ohara [4]’s $\alpha$-divergence (equivalent for $\alpha=1-2\rho$):
${D}_{\rho}^{\textnormal{Tsallis}}(q,p)$ $=-\underset{Z\sim q(z)}{\mathbb{E}}\log_{\rho}\frac{p(z)}{q(z)}=-\int q(z)^{\rho}[\log_{\rho}p(z)-\log_{\rho}q(z)]dz$ (21)
$=\frac{1}{1-\rho}[1-\int q(z)^{\rho}p(z)^{1-\rho}dz].$ (22)
Amari and Ohara [4, Eq. 40] propose a $\rho$-divergence, which is a Bregman divergence based on the convex $\rho$-cumulant function, $A_{\rho}$ between two $\rho$-exponential densities with form $\log_{\rho}q=\eta_{q}z-A_{\rho}(\eta_{q})$. Differentiating the $\rho$-cumulant $A_{\rho}$ yields a dual parametrization of the $\rho$-exponential family that is equal to the escort mean $q=\mathbb{E}_{\tilde{q}}z$, where $\tilde{q}(z;\eta)=\frac{q(z;\eta)^{\rho}}{\int q(z;\eta)^{\rho}dz}$ is the escort distribution and the escort expectation parameter $q$ is denoted by the name of the distribution. Amari and Ohara [4] show that the $\rho$-divergence is a “conformal transformation” of ${D}_{\rho}^{\textnormal{Tsallis}}$. Note also that if we multiply Amari’s $\alpha$-divergence by $\frac{\rho}{\int p_{i}^{\rho}}$ (for $\rho=\frac{1-\alpha}{2}$) we retrieve the $\rho$-divergence:
$\begin{split}D_{\rho}(q,p)&=\frac{\rho}{\int q_{i}^{\rho}}\underbrace{\frac{1}{\rho}D^{\textnormal{Tsallis}}_{\rho}(q,p)}_{\textnormal{Amari's }\alpha\textnormal{-divergence}}=\frac{-\int q_{i}(\log_{\rho}p_{i}-\log_{\rho}q_{i})}{\int q_{i}^{\rho}}=\underset{{\color[rgb]{1,0,0}\tilde{q}}}{\mathbb{E}}[\log_{\rho}q_{i}-\log_{\rho}p_{i}]\\ &=\underset{{\color[rgb]{1,0,0}\tilde{q}}}{\mathbb{E}}[\eta_{q}z-A_{\rho}(\eta_{q})-(\eta_{p}z-A_{\rho}(\eta_{p}))]=\eta_{q}q-A_{\rho}(\eta_{q})-(\eta_{p}q-A_{\rho}(\eta_{p}))\\ &=A_{\rho}(\eta_{p})-A_{\rho}(\eta_{q})-\underbrace{q}_{=\nabla A_{\rho}(\eta_{q})}(\eta_{p}-\eta_{q})=\textnormal{Bregman}_{\color[rgb]{1,0,0}{A_{\rho}}}(\eta_{p},\eta_{q})\\ &=A^{*}_{\rho}(q)+A_{\rho}(\eta_{p})-\eta_{p}q=L_{\color[rgb]{1,0,0}A^{*}_{\rho}}(\eta_{p};q).\end{split}$ (23)
For the expressions in the last line we use the Legendre duality $\eta_{q}q-A_{\rho}(\eta_{q})=A^{*}_{\rho}(q)$. $L_{\Omega}$ coincide with the Bregman divergence generated by a strictly convex and differentiable $\Omega$, when in addition $\Omega$ is a smooth Legendre type function, meaning that its gradient explodes in the boundary of its domain [10, §3.2].
The term $L_{A^{*}_{\rho}}(\eta_{p};q)$ is a divergence between distributions $p$ and $\tilde{q}$, which is different from the FY loss/divergence that we use in the main text between prior canonical parameters $\eta_{p}$ and variational mean parameters $q$ since the dual of the $\rho$-cumulant $A^{*}_{\rho}$ is different from the $\Omega$ that we use in the FY term. We discuss the difference below.

### v3

@@ 3.4 Gradient Backpropagation
We also propose a simple and versatile gradient backpropagation method to perform FYVI for general models. We use it in two of the experiments below. There, we optimize the FYELBO using amortized inference [Gershman], where the variational distribution is parametrized by an encoder network that takes $x$ as input. We denote the resulting variational distribution by $q_{\phi}(z|x)$, where $\phi$ are the inference network parameters. Instead of minimizing (21) through alternating minimization, we use end-to-end gradient backpropagation combined with a reparametrization trick to update the encoder parameters $\phi$, the decoder parameters $\theta$, and optionally any parameters associated with $\eta$ (our FYVAEs use standard fixed priors). It is also obviously possible to use a score function estimator instead of reparametrization. We also note that generalized natural-gradient algorithms could be employed here.
@@ 4 EXPERIMENTS
We illustrate FYVI through several experiments. The first is GMM-based clustering, where we recover the standard and hard versions of EM with particular choices of entropic regularizer, in addition to a new sparse EM algorithm. [^2] This sparse variant of EM is a different instance of FYVI than those proposed by dukkipati2016mixture and inoue2013q, which are standard mixtures of $(2-\rho)$-Gaussians. We also use FYVI to train $\beta$-VAEs with compact support posteriors on MNIST and FashionMNIST with both sparse and standard multivariate Bernoulli observation models. Finally, we combine a Gaussian prior with sparse observation models to model a sparse empirical distribution in a neural variational document model.
@@
Our focus is not computational scale. Computational cost depends on the particular model, entropic index $\rho$, and parametrization. Therefore, the computational cost depends on model-specific choices, rather than the generalization of standard VI to FYVI. Tradeoffs between FY and KL regularization similarly depend on model specification; there is no general comparison to be made. In our experiments, computational tradeoffs between FY regularizers and KL regularizers are negligible. There are efficient algorithms for computing sparsemax ($\rho=2$) and $1.5$-entmax; otherwise computing softmax is faster. We use reparametrization tricks in all the VAE experiments: The classical Gaussian one [Blei2017], and a $(2-\rho)$-Gaussian reparametrization [martins2022sparse]. We keep computational resources equal between compared models.
@@
Figure 3: We vary $\rho\in\{0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0\}$ and plot from left to right the resulting adjusted mutual information, adjusted Rand index, and silhouette score. In the right-most panel we quantify the E-step sparsity by counting the number of clusters assigned 0 mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned 0 proportion is three. The shaded regions represent standard errors from the mean values across 5 seeds.
4.1 FYEM for GMMs
We set $\mathcal{L}_{\theta}(x_{i},z_{i})=-\log\mathcal{N}(x_{i};\mu_{z_{i}},\Sigma_{z_{i}})$ as Gaussian negative log-likelihoods. Choosing $\Omega=\Omega_{1}$ recovers the standard EM algorithm for GMM estimation, where all the components of $\hat{q}_{i}^{\mbox{new}}$ are strictly positive. We also use other regularizers, $\Omega_{\rho}$ with $\rho>1$, which yield a new adaptively sparse EM algorithm for GMMs. In Appendix B we review the classical EM and hard EM algorithms compared in Figure 2, and give details of how the clusters are generated and evaluated in Table 1. The sparse EM algorithm interpolates between the well known standard and hard EM variants and is thereby robust to the outliers, similarly to hard EM, while still retaining soft-assignments as in the classical EM. We do not report the log marginal likelihood since this is maximized by standard EM, while sparse and hard EM optimize the FYVI objectives. Figure 3 illustrates how average sparsity and the unsupervised clustering evaluation metrics we use vary with $\rho$.
Setting $\Omega=\Omega_{\rho}$ to the Tsallis negentropy (14) with $\rho>1$, yields the new sparse version of (22):
${\hat{q}_{i}^{\mbox{new}}}=$ $\rho\mbox{-entmax}\Bigl(\frac{\pi_{1}^{\rho-1}}{\rho-1}+\log\bigl(\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1})\bigr),$...,
$\frac{\pi_{K}^{\rho-1}}{\rho-1}+\log\bigl(\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K}\bigr)\Bigr);$ (25)
see [peters2019sparse] for efficient implementations of the $\rho\mbox{-entmax}$ function for certain values of $\rho$.
When $\rho>1$ the output of $\rho\mbox{-entmax}$ can have both zero and non-zero components. The novel feature of sparse EM is the possible presence of zeros in $\hat{q}_{i}$, meaning that sample $x_{i}$ will not contribute to the update of the parameter estimates of the corresponding components. In standard EM all samples affect the parameter estimates of all the components. In hard EM, each sample only affects the parameters of the “closest” component, as in the case of $K$-means clustering (with the important difference that $K$-means clustering does not estimate component covariances).
@@ 4.1 FYEM for GMMs
Table 1: Clustering quality metrics (higher is better) for GMM-based clustering using the standard, hard, and sparse (ours) versions of EM; $\rho=2$ in sparse EM. Priors are learned according to Eq.(23). Values are averaged over 5 random seeds $\pm$ standard errors.
E-Step Adj. MI Adj. RI Silhouette
Classic .606 $\pm$ .005 .531 $\pm$ .002 .345 $\pm$ .014
Hard .537 $\pm$ .039 .348 $\pm$ .019 .207 $\pm$ .057
Sparse .636 $\pm$ .008 .476 $\pm$ .017 .393 $\pm$ .018
Table 2: VAE $\ell_{1}$ reconstruction error on MNIST and FashionMNIST datasets over 5 seeds $\pm$ std. error.
Observation Model Latent Variable Mnist FashionMnist
Product of independent Bernoulli’s, $\rho^{\prime}=1$ Gaussian, $\rho=1$ 13.236 $\pm$ .004 34.049 $\pm$ .019
Biweight, $\rho=1.5$ 12.065 $\pm$ .021 32.480 $\pm$ .018
Epanechnikov, $\rho=2$ 12.086 $\pm$ .012 32.537 $\pm$ .025
Product of $\Psi_{2}$ sparse binary distributions, $\rho^{\prime}=2$ Gaussian, $\rho=1$ 12.567 $\pm$ .017 32.881 $\pm$ .027
Biweight, $\rho=1.5$ 9.261 $\pm$ .009 25.970 $\pm$ .086
Epanechnikov, $\rho=2$ 9.144 $\pm$ .020 25.851 $\pm$ .072
@@ 4.2 Fenchel-Young $\beta$-VAE
In Kingma2014’s classic VAE a Gaussian variational posterior and a $\mathcal{N}(0,I)$ prior are combined with a product of independent Bernoulli’s as observation model, representing the pixels of MNIST digit images. The standard $\beta$-VAE baseline objective is reviewed in in Appendix C. We experiment with compact support $(2-\rho)$-Gaussian333martins2022sparse call this family $\beta$-Gaussian. variational posteriors. We also compare the baseline product of Bernoulli’s decoder with a product of binary $\Psi_{2}$ deformed exponential family distributions. The FYELBO objective is given as follows:
$\sum\limits_{i=1}^{N}\underset{z\sim q_{i}}{\mathbb{E}}\left[\ell_{\Psi_{\rho^{\prime}}}(\theta(z);x_{i})\right]+\beta D_{\Omega_{\rho}}(\eta_{\mathcal{N}(0,\textnormal{I})};q_{i}).$
@@
It is interesting to note that sparsity in the encoder and in the decoder separately contribute to improved performance. We report average $\ell_{1}$ reconstruction error as a means of comparing the different models, since the FYELBO values themselves are not comparable. In all cases we tune $\beta$ by cross-validation to prevent the posterior from collapsing to the prior.
In Appendix D we report improved $\ell_{1}$ reconstruction error in a small document model VAE, using sparsity only in the observation model. This can be seen as a generalized Bayesian treatment of sparse observation models, falling under the cases of generalized VI described in [bissiri2016general, JMLR:v24:22-0291, knoblauch2022optimization]. In preliminary experiments with sparse posteriors we did not observe improvements over the baseline NVDM of miao2016neural. This is intuitively reasonable: While the empirical distribution over words is sparse, one latent Multivariate Gaussian associated with each Multinomial (or $\Omega_{\rho}$-sparse) observation through amortized variational inference does not lead to improved reconstruction if its domain is truncated. In the image VAE data, similar patterns of pixels are 0 across images, for example on the boarders, while in the document model VAEs there is no word of the vocabulary that ought to be assigned no probability mass in all “topic dimensions”.
@@ 5 CONCLUSION
We have introduced Fenchel-Young variational inference (FYVI), expanding the class of variational posteriors to the $(2-\rho)$-deformed exponential family. FYVI exploits a general class of entropic regularizers to obtain sparse discrete and continuous posterior distributions, and distinctly extends classical Bayesian inference compared to previous generalized Bayesian and variational inference proposals like those of bissiri2016general, knoblauch2022optimization, and daudel2023monotonic. We focus on cases where the entropic index $\rho$ is greater than $1$, which leads to sparse posteriors previously unexplored in the related literature. Note that the nature of sparsity of $(2-\rho)$-exponential family posteriors is qualitatively distinct from the sparse MAP estimates that are obtained through sparsity-inducing priors. We define and lower bound the FY variational free energy (FYVFE), and describe EM-style and amortized VI algorithms for estimating arbitrary generative FYVI models. We derive and experiment with a sparse EM algorithm for Gaussian mixtures, demonstrating better clustering quality than classical and hard versions of EM on synthetic data with overlapping clusters and outliers. In our image FYVAE experiments with continuous finite support posteriors, we obtain the best performance when combining a latent $(2-\rho)$-Gaussian posterior with a sparse support observation model. We also show how to treat sparse observation models from a generalized Bayesian perspective.
@@
FYVI can be directly extended to structured latent variables by using SparseMAP [niculae2018sparsemap] posteriors; intuitively these are sparse analogues to conditional random fields. In future work we plan to extend our analysis and experiments to the non-trivial case of heavy tailed posteriors ($\rho$ < 1). Further, FYVI is a necessary step toward characterizing conjugate FY families and generalized natural gradient descent, to account for the geometry of the space of deformed exponential family distributions. Another interesting direction is to connect FYVI with variational Rényi objectives [daudel2023alpha, daudel2023monotonic]. The recently popular $\lambda$-deformation [zhang2022lambda, wong2022tsallis, guilmeau2023variational], which aims to treat Rényi and Tsallis entropies in a unified way, may provide an illuminating bridge among alternative generalized VI perspectives. Finally, connections to the recent works from online learning by alquier2021non and PAC-Bayes [alquier2016properties, alquier2024user] are also an interesting prospect.
Appendix
Appendix A RELATION BETWEEN FY EXPRESSIONS, $\alpha$- AND $q$-DIVERGENCES
In this section, we discuss how the Fenchel-Young discrepancies introduced in the main paper are related to some well-known divergences between probability distributions from the literature. This not an exhaustive review, rather we give some context about closely related and easily confused quantities. We begin by introducing the deformed $\log$ and its inverse as these functions are necessary for defining the Tsallis relative entropy, which is closely related to amari2011geometry’s $\alpha$-divergence.
$\rho$-deformation.
@@ Appendix A RELATION BETWEEN FY EXPRESSIONS, $\alpha$- AND $q$-DIVERGENCES
The Tsallis entropy is based on the $\rho$-deformed logarithmic and exponential functions for an entropic index $\rho\in\mathbb{R}\setminus\{1\}$ — this index is called $q$ by tsallis2009introduction, amari2011geometry, naudts2009q, and $\alpha$ by martins2022sparse:
$\log_{\rho}(x):=\frac{x^{1-\rho}-1}{1-\rho},\quad x>0,\quad\exp_{\rho}(x):=\left[1+(1-\rho)x\right]_{+}^{\frac{1}{1-\rho}},\quad x\in\mathbb{R}$ (26)
where $\cdot]_{+}:=\max\{\cdot,0\}$. In the limit $\rho\rightarrow 1$ these converge to $\log(x)$ and $\exp(x)$. Note that multiplication and division inside deformed logarithms is modified as follows:
@@
$D_{\Omega}$ and related divergences.
Eq. 14 in the main paper defines the Tsallis $\rho$-negentropy scaled by $\frac{1}{\rho}$ consistently with martins2022sparse: $\Omega_{\rho}(q)=-\frac{1}{\rho}S_{\rho}(q)$ where $S_{\rho}(q)=\frac{1}{1-\rho}\mathbb{E}_{q(z)}\left[q(z)^{\rho-1}-1\right]$ is the Tsallis entropy. Additionally, we note the alternative forms, using the fact that $\log_{\rho}\left(\frac{1}{q}\right)+\log_{2-\rho}q=0$,
@@
The Tsallis $\rho$-relative entropy, given by Eq. 39 in amari2011geometry is a scalar multiple of their $\alpha$-divergence (equivalent for $\alpha=1-2\rho$):
${D}_{\rho}^{\textnormal{Tsallis}}(q,p)$ $=-\underset{z\sim q(z)}{\mathbb{E}}\log_{\rho}\frac{p(z)}{q(z)}=-\int q(z)^{\rho}\left[\log_{\rho}p(z)-\log_{\rho}q(z)\right]dz$ (29)
$=\frac{1}{1-\rho}\left[1-\int q(z)^{\rho}p(z)^{1-\rho}dz\right].$ (30)
amari2011geometry use a $\rho$-divergence (see their Eq. 40), which is a Bregman divergence based on the convex $\rho$-cumulant function $A_{\rho}$, between two $\rho$-exponential densities with form $\log_{\rho}q=\eta_{q}^{\top}z-A_{\rho}(\eta_{q})$, where density $q$ is parametrized by $\eta$ and we index canonical parameter $\eta$ by the density to write the canonical $\rho$ divergences as a Bregman divergence below. Differentiating $A_{\rho}$ yields a dual parametrization of the $\rho$-exponential family that is equal to the escort mean $q=\mathbb{E}_{\tilde{q}}[z]$, where $\tilde{q}(z;\eta)=\frac{q(z;\eta)^{\rho}}{\int q(z;\eta)^{\rho}dz}$ is the escort distribution and the escort expectation parameter $q$ is denoted by the name of the distribution. amari2011geometry show that the $\rho$-divergence is a “conformal transformation” of ${D}_{\rho}^{\textnormal{Tsallis}}$. Note also that if we multiply Amari’s $\alpha$-divergence [amari2016information] by $\frac{\rho}{\int p_{i}^{\rho}}$ (for $\rho=\frac{1-\alpha}{2}$) we retrieve the $\rho$-divergence:
$\begin{split}D_{\rho}(q,p)&=\frac{\rho}{\int q_{i}^{\rho}}\underbrace{\frac{1}{\rho}D^{\textnormal{Tsallis}}_{\rho}(q,p)}_{\textnormal{Amari's }\alpha\textnormal{-divergence}}=\frac{-\int q_{i}(\log_{\rho}p_{i}-\log_{\rho}q_{i})}{\int q_{i}^{\rho}}=\underset{{\tilde{q}}}{\mathbb{E}}[\log_{\rho}q_{i}-\log_{\rho}p_{i}]\\ &=\underset{{\tilde{q}}}{\mathbb{E}}[\eta_{q}^{\top}z-A_{\rho}(\eta_{q})-(\eta_{p}^{\top}z-A_{\rho}(\eta_{p}))]=\eta_{q}^{\top}q-A_{\rho}(\eta_{q})-(\eta_{p}^{\top}q-A_{\rho}(\eta_{p}))\\ &=A_{\rho}(\eta_{p})-A_{\rho}(\eta_{q})-\underbrace{q^{\top}}_{=\nabla A_{\rho}(\eta_{q})}(\eta_{p}-\eta_{q})=\textnormal{Bregman}_{{A_{\rho}}}(\eta_{p},\eta_{q})\\ &=A^{*}_{\rho}(q)+A_{\rho}(\eta_{p})-\underset{{\tilde{q}}}{\mathbb{E}}[z^{\top}\eta_{p}].\end{split}$ (31)
For the expressions in the last line we use the Legendre duality $\eta_{q}^{\top}q-A_{\rho}(\eta_{q})=A^{*}_{\rho}(q)$.